OpenStudy (anonymous):

SOMEONE HELP OR MY SCHOOl WILL KICK ME OUT It's my semester 2 exam!! I have all my answers I just need you to verify they are correct!!!

OpenStudy (anonymous):

The function f(x)=16(5)^x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (abb0t):

You will be kicked out no matter what because you are cheating, which can result in expulsion from any academia.

OpenStudy (abb0t):

Any form of outside help that places you at a higher advantage amongst other peers on an exam is cheating.

OpenStudy (anonymous):

The answers : A. f(x) = 16(5)^x; growth rate 500% B. f(x) = 16(5)^2x; growth rate 8% C. f(x) = 16(2.24)^x; growth rate 8% D. f(x) = 16(2.24)^2x; growth rate 24%

OpenStudy (anonymous):

In other words of my illustrious friend over here...you're screweeeeeedddddddddd.